OpenStudy (anonymous):

Solve –4x + 7 = x + 2

OpenStudy (angela210793):

-4x-x=2-7 -5x=-5/-5 both sides x=1

OpenStudy (aroub):

-4x-x=-7+2 -5x=-5 x=1

OpenStudy (anonymous):

x=1